OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x - 2y = -4 3x + y = 12 -2y - 4 2y - 4 2y + 4 -2y + 4

OpenStudy (anonymous):

\(-x - 2y = -4\) Multiply by -1 both the sides: \(x + 2y = 4\) So now subtract 2y from both the sides: \(x = 4 - 2y\)

OpenStudy (anonymous):

So what choice is this ??

OpenStudy (anonymous):

@rebecca1233 I am asking to you ..

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

Really ?? -2y - 4?? 4 - 2y ?? Are they same ??

OpenStudy (anonymous):

\(x = 4 - 2y\) Tell me here 4 is positive or negative ??

OpenStudy (anonymous):

noo wait

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

4 should be positive and 2y will be negative.. What option indicates this ??

OpenStudy (anonymous):

last one ?

OpenStudy (anonymous):

Absolutely..

OpenStudy (anonymous):

Well Done..